OpenStudy (anonymous):

could someone help me pwease (: write an equation in slope intercept form for the line that satisfies the following condition slope 2/5 and passes through (10,-20)

OpenStudy (anonymous):

y=2/5x-24

OpenStudy (anonymous):

thank you so muuuch

OpenStudy (anonymous):

y=mx+b y=(2/5)x+b -20=(2/5)(10)+b b=-24 y=(2/5)x-24

OpenStudy (anonymous):

oh, and pls spare a medal if you could :) thnx!